OpenStudy (anonymous):

Raymund walks from his home to the movie house at 8 km/h and gets a ride back home at 40 km/h. If this total travelling time is half an hour, how long did it take Raymund to walk to the movie house and how far was it?

OpenStudy (anonymous):

10/3 km?

OpenStudy (anonymous):

where you get that ?

OpenStudy (anonymous):

Is it correct?

OpenStudy (anonymous):

i dont know.

OpenStudy (anonymous):

Alight so I got 3.33333 km distance and time 25 mints

OpenStudy (anonymous):

is this the answer?

OpenStudy (hoblos):

if x is the time needed to go to the movies => the time needed to go back home= 8/40 x => x+ 8/40 x = 1/2 => 48x/40 =1/2 => x= 40/96 = 0.4167h => the distance = 0.4167h * 8km/h = 3.3333 km

OpenStudy (anonymous):

thank you so much for helping me..